### v1

# Topological classification of black holes surrounded by quintessence

Abstract
In this paper, we consider a class of black hole solutions surrounded by quintessence as defects possessing topological characteristics within the thermodynamic space. By formulating a generalized off-shell free energy and calculating the winding number for each black hole branch through the residues of isolated first-order poles of characteristic function defined by this free energy, the winding number reflects the local properties of black holes. The topological number, being the sum of winding numbers for all black hole branches, reflects global properties. Utilizing the residue method, we discovered that the topological numbers for Schwarzschild, Kerr, and Kerr-AdS black holes surrounded by quintessence are consistent with those not surrounded by quintessence. Furthermore, 4D/5D Einstein-Gauss Bonnet black holes under a quintessence field belong to the same topological class as Kerr black holes. Thus both the rotation parameter and Gauss-Bonnet parameter in this background field affect the topological number of the black hole. Importantly, the winding numbers for generation and annihilation points can be calculated using the residues of second-order poles of the characteristic function.
I Introduction
Black hole thermodynamics, as a crucial avenue for exploring the deep connections between black hole properties and thermodynamic principles, originates from a groundbreaking realization: black holes possess thermodynamic quantities such as entropy and temperature  Hawking1 ; Hawking2 ; Hawking3 ; Hawking4 . Although substantial progress has been made in this field, a unified and systematic framework for characterizing the intrinsic features of various black hole systems is still lacking. In recent years, the emergence of the thermodynamic topology approach has provided a novel perspective. This method treats black holes as topological defects in thermodynamic space and defines topological invariants by analyzing the singularity structure of generalized free energy. Based on this framework, most black holes can be classified into three distinct topological classes: +1, 0, and –1, which reflect their global stability Wei2 . Due to its generality and computational efficiency, this approach has been widely applied to black holes in diverse settings, including rotating AdS black holes T4 ; T5 , phantom AdS black holes chen , Lovelock AdS black holes  T6 ; T7 , Born-Infeld black holes  T9 , as well as more complex configurations such as nonlinearly charged black holes  T10 , R-charged black holes T11 and hairy black holes  T8 . Furthermore, topology has also been extended to the study of black holes in Rényi statistics  T12 ; T13 . This topological framework opens new avenues for understanding the geometric nature of gravity. Naturally, it has been extended to black holes in the presence of dark energy fields.
Dark Matter (DM) and Dark Energy (DE) are two prominent issues in modern cosmology and astrophysics. Investigating these components facilitates understanding of the structure and evolution of the universe. Astronomical observations indicate that the universe is expanding at an accelerated rate due to DE, which involves a negative pressure driving the expansion. The introduction of DE has provided a reasonable explanation for this accelerated expansion Q1 ; Q2 ; Q3 . Statistically, DE accounts for about 73% of the total energy content of the universe, DM for about 23%, and ordinary baryonic matter only about 4%. Despite extensive astronomical observations, the fundamental nature and origin of DE remain a mystery. Various models have been proposed to explain DE, such as quintessence and the cosmological constant. Quintessence is characterized by an equation relating pressure $P$ and energy density $\rho_{q}$, expressed as $P=\omega_{q}\rho_{q}$. Depending on the value of the quintessence state parameter $\omega_{q}$, three different scenarios arise: phantom energy ($\omega_{q}<-1$), the cosmological constant ($\omega_{q}=-1$), and quintessence-like matter ($-1<\omega_{q}<-1/3$).
Black hole metrics surrounded by quintessence have long been a popular topic. Kiselev first proposed a Schwarzschild black hole solution in the presence of a quintessence field in Q4 . Following this work, many black hole solutions surrounded by typical dark energy fields have been derived. Toshmatov et al. extended Kiselev’s solution to rotating black holes using the Newman-Janis algorithm Q5 ; Q6 , as shown in Q7 . Xu et al. also applied similar methods to obtain the corresponding Kerr–Newman–AdS solution Q8 . By examining the thermodynamic consistency of these black holes, Ref. Q9 refined and further generalized these solutions. In Q10 , Benavides-Gallego et al. considered a rotating black hole with magnetic charge surrounded by quintessence and further analyzed the event horizon, ergosphere, and angular velocity of zero angular momentum observers. More importantly, they argued that dark energy significantly affects black hole dynamics. Moreover, Ndogmo et al. investigated the thermodynamic behavior of rotating black holes with nonlinear magnetic charge in a quintessence field, confirming the second law of thermodynamics and revealing phase transitions Q20 .
This work investigates the thermodynamic properties of typical black holes in such a dark energy field from a topological perspective. The structure of the paper is as follows: Section II introduces the topological number of Schwarzschild black holes with a quintessence background; Section III analyzes Kerr and Kerr-AdS black holes in this context; Section IV discusses the effect of the quintessence field on the thermodynamic topological number of EGB black holes; the final section provides our conclusions.
II Schwarzschild black hole
We begin by reviewing the work of Kiselev  Q4 , who obtained an exact solution to the Einstein field equations representing a static and spherically symmetric black hole surrounded by quintessence matter. In his formulation, the spacetime metric expresses the general form:
$ds^{2}=-N(r)dt^{2}+\frac{1}{N(r)}dr^{2}+r^{2}d\Omega_{2}^{2},$ (1)
for Schwarzschild black hole, $N(r)$ can be infored as
$N(r)=1-\frac{2M}{r}-\frac{c}{r^{3\omega_{q}+1}},$ (2)
here, $M$ denotes the black hole mass, and $c$ is a positive normalization factor linked to the quintessence energy density. When the equation-of-state parameter lies within the interval $-1<\omega_{q}<-1/3$, quintessence effectively accounts for the observed cosmic acceleration. Note that when $c=0$ the metric reduces to the standard Schwarzschild black hole metric. In this section, we focus on a typical dark energy model and adopt the representative parameter $\omega_{q}=-2/3$. For this quintessence model, the energy-momentum tensor takes the following form:
$T_{t}^{t}$ $=T_{r}^{r}=-\rho,$ (3)
$T_{\theta}^{\theta}$ $=T_{\phi}^{\phi}=\frac{\rho}{2}\left(3\omega_{q}+1\right),$
density $\rho_{q}$ is written as
$\rho_{q}=-\frac{3}{2}\frac{c\omega_{q}}{r^{3\left(\omega_{q}+1\right)}}.$ (4)
The thermodynamic quantities related to the mass, Hawking temperature and entropy of this black hole are
$M=\frac{r_{+}}{2}-\frac{c}{2}r_{+}^{-3\omega_{q}},$ (5)
$T=\frac{1}{4\pi r_{+}}(1+\frac{3c\omega_{q}}{r_{+}^{3\omega_{q}+1}}),$ (6)
$S=\pi r_{+}^{2}.$ (7)
Next, we use the residue theorem to study the topological number of the Schwarzschild black hole surrounded by a quintessence. according to eqs. (5-7), the form of the generalized Helmholtz free energy is written as Wei2
$\mathcal{F}$ $=M-\frac{S}{\tau}$ (8)
$=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}})-\frac{\pi r_{+}^{2}}{\tau},$
$\tau$ denotes an auxiliary parameter that can be interpreted as the inverse temperature parameter of the cavity surrounding the black hole. The free energy corresponds to its on-shell value when $\tau=1/T$. According to the $\phi$-mapping theory Duan , a vector field is defined as
$\phi=(\partial_{r_{+}}\mathcal{F},-\cot\Theta\csc\Theta),$ (9)
$\Theta$ is an auxiliary factor introduced for convenience in topological calculations. The value $\Theta=\pi/2$ orresponds to the zero point of the vector field $\phi$. The components of the vector field $\phi$ can be written as follows:
$\phi^{r_{+}}=\frac{1}{2}(1+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (10)
$\phi^{\theta}=-\mathrm{cot}\Theta\mathrm{csc}\Theta.$ (11)
The inverse temperature parameter can be written by using the equation $\phi^{r_{h}}=0$
$\tau=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+3c\omega_{q}}=\mathcal{G}(z).$ (12)
where $\mathcal{G}(z)$ denotes a function of the complex variable $z$. Subsequently, we construct the rational complex function $\mathcal{R}(z)$ as follows:
$\mathcal{R}(z)=\frac{1}{\tau-\mathcal{G}(z)},$ (13)
according to eq. (12) and eq. (13), we can obtian
$\mathcal{R}(z)=-\frac{z^{1+3\omega_{q}}+3c\omega_{q}}{4\pi z^{2+3\omega_{q}}-\tau(z^{1+3\omega_{q}}+3c\omega_{q})}.$ (14)
Considering the typical parameter $\omega_{q}=-2/3$, it is thus derived
$\mathcal{R}(z)=\frac{1-2cz}{(1-2cz)\tau-4\pi z}=\frac{1-2cz}{\mathcal{A}(z)}.$ (15)
The analysis of $\mathcal{A}(z)=0$ shows that there is only one positive real root, $z_{1}=\frac{\tau}{2(2\pi+c\tau)}$, According to the residue theorem, $z_{1}$ is a simple isolated pole of the function $\mathcal{R}(z)$, and the residue of $\mathcal{R}(z)$ is given by $\operatorname{Res}\mathcal{R}(z_{1})=\lim_{z\to z_{1}}(z-z_{1})\mathcal{R}(z)$. Thus, we can compute the residue of the characteristic complex function $\mathcal{R}(z)$ at the singularity $z_{1}$ as
$\operatorname{Res}\mathcal{R}(z_{1})=-\frac{1-2cz_{1}}{2(2\pi+c\tau)},$ (16)
It is worth noting that the sign of the numerator in the above expression plays an important role in our analysis, as it directly affects the sign of the winding number. According to the condition for an extremal black hole, it is straightforward to determine that $01-2cr_{+}>$ 0. Based on the definition of the winding number equation Q21
$w_{i}=\frac{\operatorname{Res}\mathcal{R}\left(z_{i}\right)}{\left|\operatorname{Res}\mathcal{R}\left(z_{i}\right)\right|}=\operatorname{Sgn}\left[\operatorname{Res}\mathcal{R}\left(z_{i}\right)\right],$ (17)
Thus, the winding number of $z_{1}$ and the total topological number of the system are given by
$w_{1}=-1,W=-1.$ (18)
By employing the framework of topological current theory, Figure 1 presents the zero point of the component $\phi^{r_{+}}$ and the corresponding unit vector field $n$. Here $r_{0}$ denotes an arbitrary length scale determined by the size of the cavity enclosing the black hole. The figure shows that the Schwarzschild black hole surrounded by quintessence possesses a single unstable branch, characterized by a topological charge of $W=-1$. This result suggests that the winding number and topological number of the Schwarzschild black hole remain unchanged despite the influence of quintessence, aligning with findings for the standard Schwarzschild case Wei2 . In other words, the quintessence field does not modify the black hole’s topological classification. Moreover, we find that the topological number obtained via the residue theorem is consistent with that derived from the topological current method. In the following, we will further examine this conclusion by presenting additional examples.
(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane
(b) The unit vector field $n$ on the $r_{+}-\Theta$ plane
Figure 1: The topological properties of the Schwarzschild black hole surrounded by quintessence, we set $cr_{0}^{2}=0.1$.
III Kerr black hole and it’s AdS case
This section investigates the topological classification of Kerr spacetime under a quintessence field, along with its extension to an AdS background. First, in the Boyer-Lindquist coordinate system, the Kerr spacetime metric in the quintessence background is given by: kbh
$ds^{2}=$ $-\left(1-\frac{2Mr+cr^{1-3\omega_{q}}}{\Sigma}\right)dt^{2}+\frac{\Sigma}{\Delta}dr^{2}$ (19)
$-2a\sin^{2}\theta\left(\frac{2Mr+cr^{1-3\omega_{q}}}{\Sigma}\right)d\phi dt+\Sigma d\theta^{2}$
$+\sin^{2}\theta\left[r^{2}+a^{2}+a^{2}\sin^{2}\theta\left(\frac{2Mr+cr^{1-3\omega_{q}}}{\Sigma}\right)\right]d\phi^{2},$
$\Delta=r^{2}-2Mr+a^{2}-cr^{1-3\omega_{q}},\Sigma=r^{2}+a^{2}cos^{2}\theta.$ (20)
where $a$ is the rotation parameter, in the absent of the quintessence field $c$, the spacetime metric (19) coincides with Kerr metric. The mass, temperature and entropy of a black hole are respectively functions of the event horizon radius $r_{+}$
$M=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}}+\frac{a^{2}}{r_{+}}),$ (21)
$T=\frac{r_{+}^{2}-a^{2}+3c\omega_{q}r_{+}^{1-3\omega_{q}}}{4\pi r_{+}(r_{+}^{2}+a^{2})},$ (22)
$S=\pi(r_{+}^{2}+a^{2}),$ (23)
Based on these thermodynamic quantities, the expression for the free energy can be derived
$\mathcal{F}=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}}+\frac{a^{2}}{r_{+}})-\frac{\pi(r_{+}^{2}+a^{2})}{\tau}.$ (24)
The components of the vector field $\phi$ can be calculated as follows
$\phi^{r_{h}}$ $=\frac{1}{2}(1-\frac{a^{2}}{r_{+}^{2}}+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (25)
$\phi^{\theta}$ $=-\mathrm{cot}\Theta\mathrm{csc}\Theta.$
When the vector component $\phi^{r_{+}}$ is zero, we can obtain
$\tau=\frac{4\pi r_{+}^{3+3\omega_{q}}}{r_{+}^{2+3\omega_{q}}-a^{2}r_{+}^{3\omega_{q}}+3c\omega_{q}r_{+}}.$ (26)
Similar to the previous section, the characteristic function in this case is defined as
$\mathcal{R}(z)=\frac{z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q}}{4\pi z^{3+3\omega_{q}}-\tau(z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q})}.$ (27)
When $\omega_{q}=-2/3$, we simplify the above equation to
$\mathcal{R}(z)=-\frac{z^{2}-2cz^{3}-a^{2}}{4\pi z^{3}-\left(z^{2}-2cz^{3}-a^{2}\right)\tau}=-\frac{z^{2}-2cz^{3}-a^{2}}{\mathcal{A}(z)}.$ (28)
$\mathcal{A}(z)$ is regarded as a polynomial function, applying the fundamental theorem of algebra, that is, any polynomial
$y(x)=m_{0}+m_{1}x+\cdots+m_{n}x^{n},\quad m_{n}\neq 0,$ (29)
can be decomposed into
$y(x)=m_{n}\left(x-z_{1}\right)\left(x-z_{2}\right)\cdots\left(x-z_{n}\right),$ (30)
where $z_{i}$ is the root of $y(x)$,which can be a complex number or a real number. Therefore, the roots of $\mathcal{A}(z)$ correspond to the poles of the function $\mathcal{R}(z)$ and the equation  (28) can be rewritten as
$\mathcal{R}(z)=-\frac{z^{2}-2cz^{3}-a^{2}}{(z-z_{1})(z-z_{2})(z-z_{3})\dots(z-z_{n})}.$ (31)
(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane, a creation point located in $\tau/r_{0}=\tau_{c}/r_{0}=3.4438$.
(b) The roots of $\mathcal{A}(z)$.
(c) The unit vector field $n$ of $\tau/r_{0}=5$.
Figure 2: The topological properties of the kerr black hole surrounded by quintessence, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
The plot of $r_{+}$ versus $\tau$ is shown in Fig. 2(a), where a stable small black hole branch and an unstable large black hole branch can be observed. At the point $\tau_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$, we find that $(d^{2}\tau/dr_{+}^{2})>0$, indicating that $\tau_{c}$ is a generation point. From Fig. 2(b), it can be seen that when $\tau>\tau_{c}$, the polynomial has one negative real root and two positive real roots
[…]
The negative root $z_{3}$ does not satisfy the off-shell condition, so only the positive real roots are considered for the topological numbers. According to the residue theorem and the definition of the winding number equation (17), we have $w_{1}=1$, $w_{2}=-1$, resulting in a total topological number $W=0$ for the Kerr black hole in the presence of quintessence. This outcome is also inferred from the vector field $n$ depicted in Fig. 2(c). In other words, the winding and topological numbers for the Kerr black hole surrounded by quintessence match those of the standard Kerr case T5 , yet they deviate from the values of the Schwarzschild black hole under the same dark energy influence. Therefore, the spacetime rotation in the quintessence framework influences its topological characteristics.
For a special case where $\tau=\tau_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$, it is found that $\mathcal{A}(z)$ has two positive roots that coincide:
[…]
Denoting this coincident point as $z_{12}$, the characteristic function (27) can be explicitly written as
$\mathcal{R}(z)=-\frac{z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q}}{(z-z_{12})^{2}(z-z_{3})}.$ (34)
Here, the obtained characteristic function $\mathcal{R}(z)$ has a second-order pole at $z_{12}$, which is a generation point. When the order $m$ of a pole $z_{i}$ satisfies $m\geq 2$, the residue of the analytic function $\mathcal{R}(z_{i})$ is given by
$\operatorname{Res}\mathcal{R}(z_{i})=\frac{1}{(m-1)!}\lim_{z\to z_{i}}\frac{d^{m-1}}{dz^{m-1}}\left[(z-z_{i})^{m}\mathcal{R}(z)\right].$ (35)
Thus, by calculating the residue of $\mathcal{R}(z)$ at the second-order pole $z_{12}$, the winding number at this special point is determined as
$w_{12}=-1.$ (36)
We observe that the winding number of this generation point differs from the results obtained using the $\phi$-mapping topological current theory and the spinodal curve calculation. As noted in Ref. Wei2 , the winding number of a generation point computed via topological current theory is 0, implying that the winding number of such points depends on the definition method. The same applies to annihilation points, as will be seen in the following discussion.
Next, we turn to the topological number of the Kerr-AdS black hole in quintessence. In Boyer-Lindquist coordinates, the spacetime of this black hole is given by kabh
$ds^{2}=$ $-\frac{\Delta_{r}}{\Omega}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right)^{2}+\frac{\Omega}{\Delta_{r}}dr^{2}$ (37)
$+\frac{\Omega}{\breve{P}}d\theta^{2}+\frac{\breve{P}\sin^{2}\theta}{\Omega}\left(adt-\frac{\left(r^{2}+a^{2}\right)}{\Xi}d\phi\right)^{2},$
where the relevant terms are defined as
$\Delta_{r}=r^{2}+a^{2}+\left(r^{2}+a^{2}\right)\left(\frac{r^{2}}{l^{2}}\right)-2mr-cr^{1-3w},$ (38)
$\Omega=r^{2}+a^{2}\cos^{2}\theta,\quad\Xi=1-\frac{a^{2}}{l^{2}},\quad\breve{P}=1-\frac{a^{2}}{l^{2}}\cos^{2}\theta.$ (39)
Under the horizon condition $\Delta_{r}=0$, the black hole mass, entropy, and Hawking temperature are
$M=\frac{m}{\Xi^{2}},$ (40)
$S=\frac{\pi(a^{2}+r_{+}^{2})}{\Xi},$ (41)
$T=\frac{\Delta r^{\prime}(r)}{4\pi(r^{2}+a^{2})}\bigg{|}{r=r_{+}}.$ (42)
Using the relation $P=\frac{3}{8\pi l^{2}}$, the generalized free energy takes the form
$\mathcal{F}=\frac{3(3+8\pi Pr_{+}^{2})(r_{+}^{2}+a^{2})-9cr_{+}^{1-3\omega q}}{6r+(3-8\pi Pa^{2})^{2}}-\frac{3\pi(r_{+}^{2}+a^{2})}{(3-8\pi Pa^{2})\tau}.$ (43)
and the components of the vector $\phi$ can be yielded as:
$\phi^{r_{+}}=$ $\frac{9(r_{+}^{2}-a^{2})+24\pi Pr_{+}^{2}(3r_{+}^{2}+a^{2})+27c\omega qr+^{1-3\omega q}}{2r+^{2}(3-8\pi Pa^{2})^{2}}$ (44)
$-\frac{6\pi r_{+}}{3-8\pi Pa^{2}\tau},$
$\phi^{\theta}=-\cot\Theta\csc\Theta,$ (45)
with
$\tau=\frac{4\pi r_{+}^{3+3\omega q}(3-8\pi Pa^{2})}{3r_{+}^{2+3\omega q}+24\pi Pr_{+}^{4+3\omega q}-3a^{2}r_{+}^{3\omega q}+8\pi Pa^{2}r_{+}^{2+3\omega_{q}}+9c\omega qr_{+}}.$ (46)
To determine the winding number, we define
$\mathcal{R}(z)=\frac{24\pi Pz^{4+3\omega_{q}}-3z^{3\omega_{q}}a^{2}+z^{2+3\omega_{q}}\left(3+8\pi Pa^{2}\right)+9cz\omega_{q}}{\tau\left(24\pi Pz^{4+3\omega_{q}}-3z^{3\omega_{q}}a^{2}+z^{2+3\omega_{q}}\left(3+8\pi Pa^{2}\right)+9cz\omega_{q}\right)+4\pi z^{3+3\omega_{q}}\left(-3+8\pi Pa^{2}\right)}.$ (47)
For $\omega_{q}=-2/3$, this simplifies to
$\mathcal{R}(z)=\frac{3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}}{4\pi z^{3}\left(-3+8P\pi a^{2}\right)+\left(3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}\right)\tau}.$ (48)
Taking the denominator of the equation, we define a polynomial function $\mathcal{A}(z)$ as
$\mathcal{A}(z)=$ $4\pi z^{3}\left(-3+8P\pi a^{2}\right)+\left(3z^{2}-6cz^{3}+24P\pi z^{4}\right.$ (49)
$\left.+8P\pi a^{2}z^{2}-3a^{2}\right)\tau.$
Figure 3(a) shows the variation of $r_{+}$ with $\tau$ where a generation point and an annihilation point (where $(d^{2}\tau/dr_{+}^{2})<0$) are marked with black dots. These points divide the black hole into three branches: a small black hole branch, an intermediate black hole branch, and a large black hole branch. To compute the topological numbers, we plot the roots of $\mathcal{A}(z)$ for three cases. For $\tau<\tau_{a}$ or $\tau>\tau_{b}$, Figure 3(b) shows the distribution of roots when $\mathcal{A}(z)=0$. In these cases, there is one positive real root and one negative real root, corresponding to a winding number and topological number of 1. Additionally, for $\tau_{a}<\tau<\tau_{b}$. In these cases, there is one positive real root and one negative real root, corresponding to a winding number and topological number of 1. Additionally, for
[…]
According to the residue theorem, the winding numbers for the positive real roots are $w_{1}=1$, $w_{2}=-1$, $w_{3}=1$, resulting in the same topological number $W=1$. This can also be seen in Fig. 3(c), which displays the unit vector field $n$ for $\tau_{a}<\tau<\tau_{b}$. Therefore, the Kerr–AdS black hole in a quintessence background exhibits a distinct topological number compared to Kerr black hole.
In Fig. 3(a), the generation and annihilation points are denoted by $\tau_{a}$ and $\tau_{b}$, respectively. The calculation of the topological number for $\tau=\tau_{a}$ using the residue theorem was discussed in the previous section and will not be repeated here. Below, we compute the topological number for $\tau=\tau_{b}$, where $z_{3}=z_{2}$
[…]
Denoting this coincident point as $z_{23}$, the characteristic function becomes
$\mathcal{R}(z)=\frac{3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}}{(z-z_{1})(z-z_{23})^{2}(z-z_{4})}.$ (52)
Using the residue theorem, we calculate the residue of the characteristic function at its second-order pole $z_{23}$, yielding the topological number for this point:
$w_{23}=1.$ (53)
(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane, a creation point and a annihilation located in $\tau/r_{0}=\tau_{a}/r_{0}=3.0101$ and $\tau/r_{0}=\tau_{b}/r_{0}=4.2354$.
(b) The roots of $\mathcal{A}(z)$.
(c) The unit vector field $n$ of $\tau/r_{0}=3.5$.
Figure 3: The topological properties of the Kerr-AdS black hole surrounded by quintessence, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
IV Einstein-Gauss-Bonnet black hole
In this section, we focus on the thermodynamic topological number associated with black hole entropy in complex forms, such as black hole solutions surrounded by dark energy in the Einstein-Gauss-Bonnet (EGB) gravity theory. The EGB action is given by:
$S=\frac{1}{16\pi}\int d^{D}x\sqrt{-g}\left[R+\alpha L_{GB}\right],$ (54)
here, $R$ the Richie tensor,$\alpha$ is the Gauss-Bonnet constant, Gauss-Bonnet Lagrange $L_{GB}$ by
$L_{GB}=R^{2}-4R_{ab}R^{ab}+R_{abcd}R^{abcd}.$ (55)
Consider the four-dimensional EGB black hole solution in the presence of quintessence matter. The corresponding spacetime metric is given by: 4DQ
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ (56)
where
$f_{\pm}=1+\frac{r^{2}}{2\alpha}(1\pm\sqrt{1+\frac{8\alpha M}{r^{3}}+\frac{8\alpha q}{r^{3\omega_{q}+3}}}).$ (57)
It is evident that when $q=0$, the solution simplifies to the four-dimensional EGB black hole proposed by Glavan and Lin  Glavan . In the limit $\alpha\to 0$, the negative branch recovers the Schwarzschild black hole immersed in quintessence matter, as described in  Q4 . Taking both $q=0$ and $\alpha\to 0$, the metric reduces to the classical Schwarzschild solution. By adjusting the state parameter $\omega_{q}$, a variety of black hole configurations can be realized. For example, when $\omega_{q}=0$, , corresponding to the absence of quintessence, the metric again reduces to the standard 4D EGB black hole  Glavan . The function $f(r)$ and the location of the event horizon are highly sensitive to the parameters $\omega_{q}$, $q$, and the Gauss–Bonnet coupling $\alpha$. To ensure the physical validity of the quintessence field, $\omega_{q}$ must remain within the interval $-1<\omega_{q}<-1/3$, , which ensures a positive energy density of the dark energy component, given by  4D
$\rho_{q}=-6w_{q}q\frac{1}{r^{3\left(1+w_{q}\right)}}.$ (58)
In equation (57), the "$\pm$" symbol represents the two branches of the solution. For large horizon radius $r$, equation (57) reduces to the four-dimensional Schwarzschild black hole solution surrounded by quintessence matter. Therefore, the negative branch of the solution is well-behaved. Consequently, we consider the negative branch solution in the following analysis, from which several thermodynamic quantities of the black hole can be derived as:
$M=\frac{r_{+}}{2}(1+\frac{\alpha}{r_{+}}-\frac{2q}{r_{+}^{3\omega_{q}+1}}),$ (59)
$T=\frac{r_{+}^{3\omega_{q}+1}\left(r_{+}^{2}-\alpha\right)+6\omega_{q}qr_{+}^{2}}{4\pi r_{+}^{3\omega_{q}+2}\left(r_{+}^{2}+2\alpha\right)}.$ (60)
Now, using the first law of thermodynamics $dM=TdS$, we obtain the expression for the black hole entropy
$S=\frac{A}{4}[1+\frac{4}{r_{+}^{2}}\mathrm{log}(r_{+})],$ (61)
Based on above thermodynamic quantities, the generalized off-shell free energy of the system can be derived
$\mathcal{F}=\frac{r_{+}}{2}(1+\frac{\alpha}{r_{+}}-\frac{2q}{r_{+}^{3\omega_{q}+1}})-\frac{\pi r_{+}^{2}(r_{+}^{2}+4\alpha\mathrm{log}(r_{+}))}{r_{+}^{2}\tau}.$ (62)
From this, the component $\phi^{r_{+}}$ of the vector field $\phi$ can be determined
$\phi^{r_{+}}=\frac{1}{2}+3q\omega_{q}r_{+}^{-1-3\omega_{q}}-\frac{2\pi(r_{+}^{2}+2\alpha)}{r_{+}\tau},$ (63)
Therefore, the inverse temperature of the 4D EGB black hole is given by
$\tau=\frac{4\pi r_{+}^{3\omega_{q}}(r_{+}^{2}+2\alpha)}{r_{+}^{1+3\omega_{q}}+6q\omega_{q}},$ (64)
In this case, the characteristic function is:
$\mathcal{R}(z)=$ $-\frac{z^{1+3\omega_{q}}+6q\omega_{q}}{4\pi z^{3\omega_{q}}(z^{2}+2\alpha)-\tau(z^{1+3\omega_{q}}+6q\omega_{q})}$ (65)
$=$ $-\frac{z^{1+3\omega_{q}}+6q\omega_{q}}{\mathcal{A}(z)}.$
Figure 4: The roots of $\mathcal{A}(z)=0$ for 4D EGB black hole. we set $\alpha=0.1$ and $q=0.1$
We analyze the roots of $\mathcal{A}(z)=0$. In order to examine the poles of the complex function $\mathcal{R}(z)$ and the corresponding winding numbers, we plot $\mathcal{A}(z)$ in Fig. 4. when $\tau_{c}<\tau$, the polynomial $\mathcal{A}(z)$ admits two positive real roots $z_{1}$ and $z_{2}$, satisfying $z_{1}<z_{2}$. Accordingly, the winding numbers are $w_{1}=1$ and $w_{2}=-1$, yielding a total topological number $W=0$. This result differs from the topological number of the Schwarzschild black hole under quintessence. Furthermore, Fig. 5 presents the zero-point diagram of $\phi^{r_{+}}$, from which it can be seen that the 4D EGB black hole features a generation point but lacks an annihilation point.
Figure 5: The zero points of $\phi^{r_{+}}$ in 4D and 5D EGB black holes are shown, where the black dots represent generation points. Let the parameters be $\alpha=0.1$ and $q=0.1$
For the case of 5D, the solution of this black hole is 5DQ
$f_{\pm}=1+\frac{r^{2}}{4\alpha}(1\pm\sqrt{1+\frac{8\alpha M}{r^{4}}+\frac{8\alpha q}{r^{4\omega_{q}+4}}}),$ (66)
Likewise, by focusing solely on the negative branch, the thermodynamic quantities associated with this spacetime can be expressed as follows
$M=r_{+}^{2}(1-\frac{q}{r_{+}^{4\omega_{q}+2}}+\frac{2\alpha}{r_{+}^{2}}),$ (67)
$S=\frac{4\pi r_{+}^{3}}{3}+16\pi\alpha r_{+},$ (68)
$T=\frac{1}{2\pi r_{+}}\left[\frac{1+\frac{2q\omega_{q}}{r_{+}^{4\omega_{q}+2}}}{1+\frac{4\alpha}{r_{+}^{2}}}\right].$ (69)
For this black hole, the off-shell free energy is given by:
$\mathcal{F}=r_{+}(1-\frac{q}{r_{+}^{4\omega_{q}}}+\frac{2\alpha}{r_{+}^{2}})-\frac{4\pi r_{+}^{3}+48\pi\alpha r_{+}}{3\tau},$ (70)
which in turn yields the following expression for the $\phi^{r_{+}}$ component of the vector field
$\phi^{r_{+}}=2r_{+}+4q\omega_{q}r_{+}^{-1-4\omega_{q}}-\frac{4\pi(r_{+}^{2}+4\alpha)}{\tau},$ (71)
and
$\tau=\frac{2\pi r_{+}^{1+4\omega_{q}}(r_{+}^{2}+4\alpha)}{r_{+}^{2+4\omega_{q}}+2q\omega_{q}}.$ (72)
The characteristic complex function in this case is
$\mathcal{R}(z)=$ $-\frac{z^{2+4\omega_{q}}+2q\omega_{q}}{2\pi z^{1+4\omega_{q}}(z^{2}+4\alpha)-\tau(z^{2+4\omega_{q}}+2q\omega_{q})}$ (73)
$=$ $-\frac{z^{2+4\omega_{q}}+2q\omega_{q}}{\mathcal{A}(z)}.$
As shown in Fig. 5, we observe that the zeros of the vector field component $\phi^{r_{+}}$ for both 4D and 5D EGB black holes exhibit similar behavior. Therefore, the results for the 5D EGB black hole can be derived from those of the 4D EGB black hole. The findings indicate that for $\tau>\tau_{c}$, the winding numbers and the topological number are $w_{1}=1$, $w_{2}=-1$, and $W=0$, respectively.
Table 1: In the table below, we summarize the topological numbers of the black holes studied in this work, along with their generation points (GP) and annihilation points (AP). We also include results reported in some references for comparison.
Black hole spacetime Topological number GP AP
Schwarzschild BH Wei2 -1 0 0
Kerr BH T4 0 1 0
Kerr-AdS BH T4 1 1 or 0 1 or 0
Schwarzschild BH surrounded by quintessence -1 0 0
Kerr BH surrounded by quintessence 0 1 0
Kerr-AdS BH surrounded by quintessence 1 1 or 0 1 or 0
4D EGB BH surrounded by quintessence 0 1 0
5D EGB BH surrounded by quintessence 0 1 0
V Conclusions
This article explores the thermodynamic topological properties of black holes by utilizing the residue theorem. In this study, black hole solutions are regarded as defects in thermodynamic space, and a characteristic complex function is defined based on the generalized off-shell free energy. The singularities of this complex function are determined by the roots of the polynomial $\mathcal{A}(z)$. The winding number of each black hole branch is given by the residue of the characteristic function at its simple poles, which can be calculated using formula (17). The winding numbers reveal the local topological properties of the black hole system. By summing the winding numbers of all singular points, the global topological number is obtained, which aids in classifying different black hole spacetimes. Moreover, the winding numbers of generation points and annihilation points can be identified by the residues of the characteristic function at their second-order isolated singularities.
We studied the winding numbers and topological numbers of a family of black holes surrounded by quintessence, with the results summarized in Table I, demonstrating that the results obtained via the residue method agree with those calculated through the topological current theory. Within the chosen black hole models, we find that the Schwarzschild black hole in quintessence has a topological number of -1, with no generation or annihilation points. In contrast, the Kerr black hole and the 4D/5D Einstein-Gauss-Bonnet black holes under quintessence background have a topological number of 0, each featuring one generation point but no annihilation point. For the Kerr-AdS black hole in quintessence, its topological number matches that of the Kerr-AdS black hole without quintessence, with one generation point and one annihilation point identified. These results indicate that both the rotation parameter and the Gauss-Bonnet coupling parameter under the quintessence background alter the topological class of black holes.
Acknowledgments
The research was partially supported by the National Natural Science Foundation of China (Grant No.12265007) and the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07, QJJ-[2022]-314). Also, the research of H.H. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by project PID2023-148409NB-I00, funded by MICIU/AEI/10.13039/501100011033. Financial support of the Department of Education of the Junta de Castilla y Leon and FEDER Funds is also gratefully acknowledged (Reference: CLU-2023-1-05). Additionally, H. H. is grateful to Excellence project FoS UHK 2203/2025-2026 for the financial support.

### v2

@@
-Topological classification of black holes surrounded by quintessence
+Universal topological classes of black holes surrounded by quintessence
 Abstract
-In this paper, we consider a class of black hole solutions surrounded by quintessence as defects possessing topological characteristics within the thermodynamic space.
-By formulating a generalized off-shell free energy and calculating the winding number for each black hole branch through the residues of isolated first-order poles of characteristic function defined by this free energy, the winding number reflects the local properties of black holes.
-The topological number, being the sum of winding numbers for all black hole branches, reflects global properties.
-Utilizing the residue method, we discovered that the topological numbers for Schwarzschild, Kerr, and Kerr-AdS black holes surrounded by quintessence are consistent with those not surrounded by quintessence.
-Furthermore, 4D/5D Einstein-Gauss Bonnet black holes under a quintessence field belong to the same topological class as Kerr black holes.
-Thus both the rotation parameter and Gauss-Bonnet parameter in this background field affect the topological number of the black hole.
-Importantly, the winding numbers for generation and annihilation points can be calculated using the residues of second-order poles of the characteristic function.
+This work explores the universal topological classes of various black hole configurations immersed in a quintessence field.
+The results indicate that the Schwarzschild black hole surrounded by quintessence is of the $W^{1-}$ type, exhibiting thermodynamic instability in both extremal temperature limits.
+The introduction of rotation alters its topological nature to the $W^{0+}$ class, where the Kerr black hole displays coexistence of a stable small black hole and an unstable large black hole at low temperatures.
+In comparison, anti de Sitter (adS) black holes display different thermodynamic topologies.
+The Schwarzschild–adS solution corresponds to the $W^{0-}$ class, where an unstable small branch and a stable large branch emerge in the high temperatures.
+Meanwhile, the Kerr–adS classified as $W^{1+}$, maintains stability in both high- and low-temperature limits.
+Furthermore, the 4D/5D quintessential black hole belongs to the $W^{0+}$ class within the Einstein–Gauss–Bonnet framework.
+Overall, parameters such as rotation and Gauss-Bonnet in quintessence background field affect the topological classifications of the black holes, while the quintessence field primarily modifies the stability range without changing the topological class itself.
 I Introduction
-Black hole thermodynamics, as a crucial avenue for exploring the deep connections between black hole properties and thermodynamic principles, originates from a groundbreaking realization: black holes possess thermodynamic quantities such as entropy and temperature  Hawking1 ; Hawking2 ; Hawking3 ; Hawking4 .
+Black hole thermodynamics, as a crucial avenue for exploring the deep connections between black hole properties and thermodynamic principles, originates from a groundbreaking realization: black holes possess thermodynamic quantities such as entropy and temperature H1 ; H2 ; H3 ; H4 .
 Although substantial progress has been made in this field, a unified and systematic framework for characterizing the intrinsic features of various black hole systems is still lacking.
@@
 Furthermore, topology has also been extended to the study of black holes in Rényi statistics T12 ; T13 .
+Recently, Ref.
+T14 further refined the topological classification of black hole thermodynamics into four universal topological types.
+Subsequently, Ref.
+T15 expanded the existing classification framework.
+The new topological classification method has already been applied to other types of black holes T16 ; T17 ; T18 ; T19 .
 This topological framework opens new avenues for understanding the geometric nature of gravity.
@@
 Quintessence is characterized by an equation relating pressure $P$ and energy density $\rho_{q}$, expressed as $P=\omega_{q}\rho_{q}$.
-Depending on the value of the quintessence state parameter $\omega_{q}$, three different scenarios arise: phantom energy ($\omega_{q}<-1$), the cosmological constant ($\omega_{q}=-1$), and quintessence-like matter ($-1<\omega_{q}<-1/3$).
+Depending on the value of the quintessence state parameter $\omega_{q}$, three different scenarios arise: phantom energy ($\omega_{q}<-1$), the cosmological constant ($\omega_{q}=-1$), and quintessence-like matter ($1<\omega_{q}<-1/3$).
 Black hole metrics surrounded by quintessence have long been a popular topic.
@@
 investigated the thermodynamic behavior of rotating black holes with nonlinear magnetic charge in a quintessence field, confirming the second law of thermodynamics and revealing phase transitions Q20 .
-This work investigates the thermodynamic properties of typical black holes in such a dark energy field from a topological perspective.
-The structure of the paper is as follows: Section II introduces the topological number of Schwarzschild black holes with a quintessence background; Section III analyzes Kerr and Kerr-AdS black holes in this context; Section IV discusses the effect of the quintessence field on the thermodynamic topological number of EGB black holes; the final section provides our conclusions.
-II Schwarzschild black hole
+The structure of this paper is organized as follows: in the next section, we investigate the topological class of Schwarzschild black hole surrounded by quintessence.
+Section III shifts the focus to the thermodynamic topological classification of Kerr black hole in the presence of quintessence.
+Sections IV and V examine the Schwarzschild–AdS and Kerr–AdS black holes within the quintessence matter, respectively.
+Section VI discusses the topological characteristics of Einstein–Gauss–Bonnet (EGB) black holes.
+The final section provides our conclusions.
+II Schwarzschild black hole surrounded by quintessence
 We begin by reviewing the work of Kiselev Q4 , who obtained an exact solution to the Einstein field equations representing a static and spherically symmetric black hole surrounded by quintessence matter.
 In his formulation, the spacetime metric expresses the general form:
-$ds^{2}=-N(r)dt^{2}+\frac{1}{N(r)}dr^{2}+r^{2}d\Omega_{2}^{2},$ (1)
+$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}d\Omega_{2}^{2},$ (1)
 for Schwarzschild black hole, $N(r)$ can be infored as
-$N(r)=1-\frac{2M}{r}-\frac{c}{r^{3\omega_{q}+1}},$ (2)
+$f(r)=1-\frac{2M}{r}-\frac{c}{r^{3\omega_{q}+1}},$ (2)
 here, $M$ denotes the black hole mass, and $c$ is a positive normalization factor linked to the quintessence energy density.
-When the equation-of-state parameter lies within the interval $-1<\omega_{q}<-1/3$, quintessence effectively accounts for the observed cosmic acceleration.
+When the equation-of-state parameter lies within the interval $1<\omega_{q}<-1/3$, quintessence effectively accounts for the observed cosmic acceleration.
 Note that when $c=0$ the metric reduces to the standard Schwarzschild black hole metric.
-In this section, we focus on a typical dark energy model and adopt the representative parameter $\omega_{q}=-2/3$.
-For this quintessence model, the energy-momentum tensor takes the following form:
+In this paper, we focus on a typical dark energy model and adopt the representative parameter $\omega_{q}=-2/3$.
+For this quintessence model, the energy-momentum tensor takes the following form
 $T_{t}^{t}$ $=T_{r}^{r}=-\rho,$ (3)
@@
 $S=\pi r_{+}^{2}.$ (7)
-Next, we use the residue theorem to study the topological number of the Schwarzschild black hole surrounded by a quintessence.
+It is straightforward to confirm that, for the Schwarzschild black hole surrounded by quintessence (Sch–BH-$\omega_{q}$), the Hawking temperature tends to vanish as the event horizon radius $r_{+}$ approaches infinity, whereas it becomes divergent when $r_{+}$ approaches the minimum horizon radius $r_{m}$.
+As a result, the inverse temperature $\beta(r_{+})$ shows
+$\beta(r_{m})=0,\beta(r_{\infty})=\infty.$ (8)
+At above asymptotic limits, defect curve $\beta(r_{+})$ is analytical in the range $(r_{m},\infty)$.
+The topological properties of the Sch–BH-$\omega_{q}$ are fundamentally rooted in the concept of generalized off shell free energy.
 According to eqs.
-(5-7), the form of the generalized Helmholtz free energy is written as Wei2
-$\mathcal{F}$ $=M-\frac{S}{\tau}$ (8)
-$=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}})-\frac{\pi r_{+}^{2}}{\tau},$
+(5-7), the form of the free energy is written as Wei2
+$\mathcal{F}$ $=M-\frac{S}{\tau}=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}})-\frac{\pi r_{+}^{2}}{\tau},$ (9)
 $\tau$ denotes an auxiliary parameter that can be interpreted as the inverse temperature parameter of the cavity surrounding the black hole.
-The free energy corresponds to its on-shell value when $\tau=1/T$.
-According to the $\phi$-mapping theory Duan , a vector field is defined as
-$\phi=(\partial_{r_{+}}\mathcal{F},-\cot\Theta\csc\Theta),$ (9)
+The free energy corresponds to its on-shell value when $\tau=\beta=1/T$.
+According to Duan’s $\phi$-mapping theory Duan1 ; Duan2 , we introduce an additional parameter $\Theta\in(0,\pi)$, accordingly, the two-component vector $\phi$ is specified by the gradient of as $\tilde{\mathcal{F}}=\mathcal{F}+\frac{1}{\sin\Theta}$ T14
+[…]
 $\Theta$ is an auxiliary factor introduced for convenience in topological calculations.
 The value $\Theta=\pi/2$ orresponds to the zero point of the vector field $\phi$.
-The components of the vector field $\phi$ can be written as follows:
+Based on $\phi$-mapping theory, each black hole state is assigned a winding number reflecting its topological charge.
+Locally stable states correspond to $w=+1$, while unstable ones have w=1w=-1.
+The sum of all local winding numbers defines the global topological number $W$, which determines the black hole’s overall topological class.
+We establish the components of the vector field $\phi$ as:
 $\phi^{r_{+}}=\frac{1}{2}(1+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (11)
@@
 The inverse temperature parameter can be written by using the equation $\phi^{r_{h}}=0$
-$\tau=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+3c\omega_{q}}=\mathcal{G}(z).$ (12)
-where $\mathcal{G}(z)$ denotes a function of the complex variable $z$. Subsequently, we construct the rational complex function $\mathcal{R}(z)$ as follows:
-$\mathcal{R}(z)=\frac{1}{\tau-\mathcal{G}(z)},$ (13)
-according to eq. (12) and eq. (13), we can obtian
-$\mathcal{R}(z)=-\frac{z^{1+3\omega_{q}}+3c\omega_{q}}{4\pi z^{2+3\omega_{q}}-\tau(z^{1+3\omega_{q}}+3c\omega_{q})}.$ (14)
-Considering the typical parameter $\omega_{q}=-2/3$, it is thus derived
-$\mathcal{R}(z)=\frac{1-2cz}{(1-2cz)\tau-4\pi z}=\frac{1-2cz}{\mathcal{A}(z)}.$ (15)
-The analysis of $\mathcal{A}(z)=0$ shows that there is only one positive real root, $z_{1}=\frac{\tau}{2(2\pi+c\tau)}$, According to the residue theorem, $z_{1}$ is a simple isolated pole of the function $\mathcal{R}(z)$, and the residue of $\mathcal{R}(z)$ is given by $\operatorname{Res}\mathcal{R}(z_{1})=\lim_{z\to z_{1}}(z-z_{1})\mathcal{R}(z)$.
-Thus, we can compute the residue of the characteristic complex function $\mathcal{R}(z)$ at the singularity $z_{1}$ as
-$\operatorname{Res}\mathcal{R}(z_{1})=-\frac{1-2cz_{1}}{2(2\pi+c\tau)},$ (16)
-It is worth noting that the sign of the numerator in the above expression plays an important role in our analysis, as it directly affects the sign of the winding number.
-According to the condition for an extremal black hole, it is straightforward to determine that $01-2cr_{+}>$ 0.
-Based on the definition of the winding number equation Q21
-$w_{i}=\frac{\operatorname{Res}\mathcal{R}\left(z_{i}\right)}{\left|\operatorname{Res}\mathcal{R}\left(z_{i}\right)\right|}=\operatorname{Sgn}\left[\operatorname{Res}\mathcal{R}\left(z_{i}\right)\right],$ (17)
-Thus, the winding number of $z_{1}$ and the total topological number of the system are given by
-$w_{1}=-1,W=-1.$ (18)
-By employing the framework of topological current theory, Figure 1 presents the zero point of the component $\phi^{r_{+}}$ and the corresponding unit vector field $n$.
-Here $r_{0}$ denotes an arbitrary length scale determined by the size of the cavity enclosing the black hole.
-The figure shows that the Schwarzschild black hole surrounded by quintessence possesses a single unstable branch, characterized by a topological charge of $W=-1$.
-This result suggests that the winding number and topological number of the Schwarzschild black hole remain unchanged despite the influence of quintessence, aligning with findings for the standard Schwarzschild case Wei2 .
-In other words, the quintessence field does not modify the black hole’s topological classification.
-Moreover, we find that the topological number obtained via the residue theorem is consistent with that derived from the topological current method.
-In the following, we will further examine this conclusion by presenting additional examples.
-(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane
-(b) The unit vector field $n$ on the $r_{+}-\Theta$ plane
-Figure 1: The topological properties of the Schwarzschild black hole surrounded by quintessence, we set $cr_{0}^{2}=0.1$.
-III Kerr black hole and it’s AdS case
-This section investigates the topological classification of Kerr spacetime under a quintessence field, along with its extension to an AdS background.
+$\tau=\beta=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+3c\omega_{q}}.$ (13)
+Figure 1: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for Sch-BH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
+Figure 1 presents the component $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane, here $r_{0}$ denotes an arbitrary length scale determined by the size of the cavity enclosing the black hole.
+The figure shows that the Sch–BH-$\omega_{q}$ possesses a single unstable branch.
+We proceed to examine the variation of the vector field $\phi$ (see Fig.
+2) along the boundary defined by Eq.
+(8).
+The complete boundary is outlined by the closed contour $C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$, whose segments and overall shape are shown in Fig.
+3, encompassing the entire parameter domain.
+Due to the construction of $\phi$, it remains perpendicular to $I_{2}$ and $I_{4}$ T14 , implying that the key asymptotic properties are determined by its behavior on $I_{1}$ and $I_{3}$.
+Specifically, as $r_{+}\to r_{m}$, the vector field directs rightward with an inclination governed by $\phi^{\Theta}$; whereas in the limit $r_{+}\to\infty$, it turns leftward, as illustrated in Fig.
+2 and Fig.
+3.
+Table I presents the orientations of the vector field $\phi$ along segments $I_{1}$ and $I_{3}$.
+[…]
+Figure 2: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Sch-BH-$\omega_{q}$, we set $\beta/r_{0}=20$.
+Figure 3: The asymptotic behavior of the $n$-vector field $\phi$ at the boundary ($C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$).
+Figure 4: Contour $\Phi_{1}$ depicts the variation of the vector field $\phi$ components along the pat $C_{1}$ show in Fig. 2 for the Sch-BH-$\omega_{q}$ case.
+Table 1: The orientation is indicated by the arrow of $\phi^{r_{+}}$ for the Sch-BH-$\omega_{q}$ and Kerr-BH-$\omega_{q}$ and the topological numbers.
+BH spacetime $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
+Sch-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\to$ $\downarrow$ -1
+Kerr-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\leftarrow$ $\downarrow$ 0
+[…]
+III Kerr black hole surrounded by quintessence
+This section investigates the topological classification of Kerr spacetime under a quintessence field.
 First, in the Boyer-Lindquist coordinate system, the Kerr spacetime metric in the quintessence background is given by: kbh
@@
 $S=\pi(r_{+}^{2}+a^{2}),$ (18)
-Based on these thermodynamic quantities, the expression for the free energy can be derived
+the asymptotic behavior of the inverse temperature $\beta(r_{+})$ for the Kerr black hole surrounded by quintessence (Kerr-BH-$\omega_{q}$) has the limits
+$\beta(r_{m})=\infty,\beta(r_{\infty})=\infty.$ (19)
+Based on above thermodynamic quantities, the expression for the free energy can be derived
 $\mathcal{F}=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}}+\frac{a^{2}}{r_{+}})-\frac{\pi(r_{+}^{2}+a^{2})}{\tau}.$ (20)
 The components of the vector field $\phi$ can be calculated as follows
-$\phi^{r_{h}}$ $=\frac{1}{2}(1-\frac{a^{2}}{r_{+}^{2}}+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (25)
+$\phi^{r_{+}}$ $=\frac{1}{2}(1-\frac{a^{2}}{r_{+}^{2}}+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (21)
 $\phi^{\theta}$ $=-\mathrm{cot}\Theta\mathrm{csc}\Theta.$
 When the vector component $\phi^{r_{+}}$ is zero, we can obtain
-$\tau=\frac{4\pi r_{+}^{3+3\omega_{q}}}{r_{+}^{2+3\omega_{q}}-a^{2}r_{+}^{3\omega_{q}}+3c\omega_{q}r_{+}}.$ (26)
-Similar to the previous section, the characteristic function in this case is defined as
-$\mathcal{R}(z)=\frac{z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q}}{4\pi z^{3+3\omega_{q}}-\tau(z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q})}.$ (27)
-When $\omega_{q}=-2/3$, we simplify the above equation to
-$\mathcal{R}(z)=-\frac{z^{2}-2cz^{3}-a^{2}}{4\pi z^{3}-\left(z^{2}-2cz^{3}-a^{2}\right)\tau}=-\frac{z^{2}-2cz^{3}-a^{2}}{\mathcal{A}(z)}.$ (28)
-$\mathcal{A}(z)$ is regarded as a polynomial function, applying the fundamental theorem of algebra, that is, any polynomial
-$y(x)=m_{0}+m_{1}x+\cdots+m_{n}x^{n},\quad m_{n}\neq 0,$ (29)
-can be decomposed into
-$y(x)=m_{n}\left(x-z_{1}\right)\left(x-z_{2}\right)\cdots\left(x-z_{n}\right),$ (30)
-where $z_{i}$ is the root of $y(x)$,which can be a complex number or a real number.
-Therefore, the roots of $\mathcal{A}(z)$ correspond to the poles of the function $\mathcal{R}(z)$ and the equation  (28) can be rewritten as
-$\mathcal{R}(z)=-\frac{z^{2}-2cz^{3}-a^{2}}{(z-z_{1})(z-z_{2})(z-z_{3})\dots(z-z_{n})}.$ (31)
-(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane, a creation point located in $\tau/r_{0}=\tau_{c}/r_{0}=3.4438$.
-(b) The roots of $\mathcal{A}(z)$.
-(c) The unit vector field $n$ of $\tau/r_{0}=5$.
-Figure 2: The topological properties of the kerr black hole surrounded by quintessence, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
-The plot of $r_{+}$ versus $\tau$ is shown in Fig.
-2(a), where a stable small black hole branch and an unstable large black hole branch can be observed.
-At the point $\tau_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$, we find that $(d^{2}\tau/dr_{+}^{2})>0$, indicating that $\tau_{c}$ is a generation point.
-From Fig.
-2(b), it can be seen that when $\tau>\tau_{c}$, the polynomial has one negative real root and two positive real roots
+$\tau=\beta=\frac{4\pi r_{+}^{3+3\omega_{q}}}{r_{+}^{2+3\omega_{q}}-a^{2}r_{+}^{3\omega_{q}}+3c\omega_{q}r_{+}}.$ (22)
+Figure 5: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for Kerr-BH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
+Figure 6: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Kerr-BH-$\omega_{q}$ for $\beta/r_{0}=5$.
+Figure 7: Contours $\Phi_{2}$ and $\Phi_{3}$ depict the variation of the vector field $\phi$ components along the pats $C_{2}$ and $C_{3}$ show in Fig. 6 for the Kerr-BH-$\omega_{q}$ case.
+The plot of $r_{+}$ versus $\beta$ is shown in Fig.
+5, where a small black hole branch and a large black hole branch can be observed.
+A special point $\beta_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$ for which $d\beta/dr_{+}=0$ is degenerate, at this point we find that $(d^{2}\beta/dr_{+}^{2})>0$, indicating that $\beta_{a}$ is a generation point.
+When we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$, the generation point located in $\beta/r_{0}=\beta_{c}/r_{0}=3.4438$.
+The characteristics of the vector field $n$ shown in Fig.
+6, together with its asymptotic behavior along the boundary segments $I_{i}$ (analogous to the case of the Sch-BH-$\omega_{q}$ illustrated in Fig.
+3), are summarized in Table I.
+It is found that in the limits $r_{h}\to r_{m}$ or $r_{h}\to\infty$, the vector field plot of $\phi$ points leftward.
+Figure 7 illustrates the variation pattern of the vector field $\phi$ along the chosen contours.
+It can be observed that the closed trajectory $\Phi_{2}$ rotates counterclockwise, leading to winding numbers of +1 and -1 for the first and second zero points, respectively.
+This result indcates a total topological number $W=0$ for the Kerr-BH-$\omega_{q}$.
+For the universal thermodynamic behavior, when $\beta\to\infty$, both a stable small black hole and an unstable large black hole coexist.
+In contrast, as $\beta\to 0$, the black hole phase disappears entirely.
+Therefore, the Kerr-BH-$\omega_{q}$ is identified with the topological class $W^{0+}$.
+Moreover, the winding numbers and universal topological class for the Kerr-BH-$\omega_{q}$ match those of the standard Kerr case T17 , yet they deviate from the values of the Schwarzschild black hole under the same dark energy influence.
+Therefore, the spacetime rotation in the quintessence framework influences its topological characteristics.
+IV Schwarzschild-adS black hole surrounded by quintessence
+Then using the ideas of Kiselev who proposed the solution of the Schwarzschild–anti de Sitter black hole surrounded by quintessence (Sch-aBH-$\omega_{q}$) Q4
+$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2},$ (23)
+with
+$f(r)=1-\frac{2M}{r}-\frac{\Lambda r^{2}}{3}-\frac{c}{r^{3\omega_{q}+1}},$ (24)
+where $\Lambda$ is the cosmological constant. Then by using Eq. (24), one gets
+$M=\frac{r_{+}}{2}-\frac{cr_{+}^{-3\omega_{q}}}{2}-\frac{\Lambda r_{+}^{3}}{6}.$ (25)
+The Hawking temperature and entropy associated with the black hole horizon can be determined
+$T=\frac{1}{4\pi r_{+}}-\frac{\Lambda r_{+}}{4\pi}+\frac{3c\omega_{q}r_{+}^{-2-3\omega_{q}}}{4\pi},$ (26)
+$S=\pi r_{+}^{2}.$
+The asymptotic behavior of the inverse temperature $\beta(r_{+})$ exhibits
+$\beta(r_{m})=0,\penalty 10000\ \beta(r_{\infty})=0.$ (27)
+The generalized off-shell Helmholtz free energy can be easily obtained
+$\mathcal{F}=\frac{r_{+}}{2}-\frac{cr_{+}^{-3\omega_{q}}}{2}-\frac{\Lambda r_{+}^{3}}{6}-\frac{\pi r_{+}^{2}}{\tau},$ (28)
+the components of the vector $\phi$ can be computed as
+$\phi^{r_{+}}=\frac{1}{2}+4\pi Pr_{+}^{2}+\frac{3c\omega_{q}r_{+}^{-1-3\omega_{q}}}{2}-\frac{2\pi r_{+}}{\tau},$ (29)
+$\phi^{\theta}=-\cot\Theta\csc\Theta,$ (30)
+and the inverse temperature is
+$\beta=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+8\pi Pr_{+}^{3+3\omega_{q}}+3c\omega_{q}}.$ (31)
+Figure 8: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for the Sch-aBH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$, and $Pr_{0}^{2}=0.1$.
+Figure 9: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Sch-aBH-$\omega_{q}$, we set $\beta/r_{0}=3$.
+Figure 10: Contours $\Phi_{4}$ and $\Phi_{5}$ depict the variation of the vector field $\phi$ components along the pats $C_{4}$ and $C_{5}$ show in Fig. 9 for the Sch-aBH-$\omega_{q}$ case.
+In this framework, the pressure term is defined by $P=-\Lambda/8\pi$.
+The plot of $r_{+}$ versus $\beta$ is shown in Fig.
+8.
+At the point $\beta_{c}=\frac{2\pi}{3\sqrt{2\pi P}-c}$ which is a degenerate point, we find that $(d^{2}\beta/dr_{+}^{2})<0$, indicating that $\beta_{c}$ is an annihilation point.
+When $\beta<\beta_{c}$, two black hole phases coexist a small branch and a large branch.
+Once $\beta>\beta_{c}$, these phases disappear, leaving no viable black hole solution.
+The annihilation point locates at $\beta_{c}/r_{0}=4.23016$ for $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
+The corresponding vector field $n$ associated with the Sch-aBH-$\omega_{q}$ is illustrated in Fig.
+9.
+As shown in the figure, this black hole exhibits two distinct zero points (depicted as black dots), which is in contrast to the Sch-BH-$\omega_{q}$ that possesses only one.
+The asymptotic orientation of the vector field $\phi$ at the boundaries $I_{i}$, as listed in Table II, reveals that as $r_{+}\to r_{m}$ or $r\to\infty$, the field tends to point rightward.
+This directional behavior indicates that the thermodynamic topology of the system corresponds to a particular universal class.
+To further elucidate the topological structure, Fig.
+10 presents the evolution of the vector field components ($\phi^{r_{+}},\phi^{\Theta}$) traced along the closed contours $C_{i}$ depicted in Fig.
+9.
+The zero point of $\Phi_{i}$ is located at the origin.
+Contours with the first and second zeros possess winding numbers -1 and +1, respectively.
 […]
-The negative root $z_{3}$ does not satisfy the off-shell condition, so only the positive real roots are considered for the topological numbers.
-According to the residue theorem and the definition of the winding number equation (17), we have $w_{1}=1$, $w_{2}=-1$, resulting in a total topological number $W=0$ for the Kerr black hole in the presence of quintessence.
-This outcome is also inferred from the vector field $n$ depicted in Fig.
-2(c).
-In other words, the winding and topological numbers for the Kerr black hole surrounded by quintessence match those of the standard Kerr case T5 , yet they deviate from the values of the Schwarzschild black hole under the same dark energy influence.
-Therefore, the spacetime rotation in the quintessence framework influences its topological characteristics.
-For a special case where $\tau=\tau_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$, it is found that $\mathcal{A}(z)$ has two positive roots that coincide:
-[…]
-Denoting this coincident point as $z_{12}$, the characteristic function (27) can be explicitly written as
-$\mathcal{R}(z)=-\frac{z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q}}{(z-z_{12})^{2}(z-z_{3})}.$ (34)
-Here, the obtained characteristic function $\mathcal{R}(z)$ has a second-order pole at $z_{12}$, which is a generation point.
-When the order $m$ of a pole $z_{i}$ satisfies $m\geq 2$, the residue of the analytic function $\mathcal{R}(z_{i})$ is given by
-$\operatorname{Res}\mathcal{R}(z_{i})=\frac{1}{(m-1)!}\lim_{z\to z_{i}}\frac{d^{m-1}}{dz^{m-1}}\left[(z-z_{i})^{m}\mathcal{R}(z)\right].$ (35)
-Thus, by calculating the residue of $\mathcal{R}(z)$ at the second-order pole $z_{12}$, the winding number at this special point is determined as
-$w_{12}=-1.$ (36)
-We observe that the winding number of this generation point differs from the results obtained using the $\phi$-mapping topological current theory and the spinodal curve calculation.
-As noted in Ref.
-Wei2 , the winding number of a generation point computed via topological current theory is 0, implying that the winding number of such points depends on the definition method.
-The same applies to annihilation points, as will be seen in the following discussion.
-Next, we turn to the topological number of the Kerr-AdS black hole in quintessence. In Boyer-Lindquist coordinates, the spacetime of this black hole is given by kabh
+Table 2: The orientation is indicated by the arrow of $\phi^{r_{+}}$ for the Sch-aBH-$\omega_{q}$ and Kerr-aBH-$\omega_{q}$ and the topological numbers..
+BH spacetime $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
+Sch-aBH-$\omega_{q}$ $\to$ $\uparrow$ $\to$ $\downarrow$ 0
+Kerr-aBH-$\omega_{q}$ $\to$ $\uparrow$ $\leftarrow$ $\downarrow$ 1
+V Kerr-adS black hole surrounded by quintessence
+Next, we turn to the topological class of the Kerr-anti de Sitter black hole surrounded by quintessence (Kerr-aBH-$\omega_{q}$).
+In Boyer-Lindquist coordinates, the spacetime of this black hole is given by kabh
 $ds^{2}=$ $-\frac{\Delta_{r}}{\Omega}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right)^{2}+\frac{\Omega}{\Delta_{r}}dr^{2}$ (32)
@@
 $S=\frac{\pi(a^{2}+r_{+}^{2})}{\Xi},$ (36)
-$T=\frac{\Delta r^{\prime}(r)}{4\pi(r^{2}+a^{2})}\bigg{|}{r=r_{+}}.$ (42)
+$T=\frac{\Delta r^{\prime}(r)}{4\pi(r^{2}+a^{2})}\bigg|_{{r=r_{+}}}.$ (37)
 Using the relation $P=\frac{3}{8\pi l^{2}}$, the generalized free energy takes the form
-$\mathcal{F}=\frac{3(3+8\pi Pr_{+}^{2})(r_{+}^{2}+a^{2})-9cr_{+}^{1-3\omega q}}{6r+(3-8\pi Pa^{2})^{2}}-\frac{3\pi(r_{+}^{2}+a^{2})}{(3-8\pi Pa^{2})\tau}.$ (43)
+$\mathcal{F}=$ $\frac{3(3+8\pi Pr_{+}^{2})(r_{+}^{2}+a^{2})-9cr_{+}^{1-3\omega q}}{6r_{+}(3-8\pi Pa^{2})^{2}}$ (38)
+$-\frac{3\pi(r_{+}^{2}+a^{2})}{(3-8\pi Pa^{2})\tau}.$
 and the components of the vector $\phi$ can be yielded as:
-$\phi^{r_{+}}=$ $\frac{9(r_{+}^{2}-a^{2})+24\pi Pr_{+}^{2}(3r_{+}^{2}+a^{2})+27c\omega qr+^{1-3\omega q}}{2r+^{2}(3-8\pi Pa^{2})^{2}}$ (44)
+$\phi^{r_{+}}=$ $\frac{9(r_{+}^{2}-a^{2})+24\pi Pr_{+}^{2}(3r_{+}^{2}+a^{2})+27c\omega_{q}r_{+}^{1-3\omega_{q}}}{2r_{+}^{2}(3-8\pi Pa^{2})^{2}}$ (39)
 $-\frac{6\pi r_{+}}{3-8\pi Pa^{2}\tau},$
@@
 with
-$\tau=\frac{4\pi r_{+}^{3+3\omega q}(3-8\pi Pa^{2})}{3r_{+}^{2+3\omega q}+24\pi Pr_{+}^{4+3\omega q}-3a^{2}r_{+}^{3\omega q}+8\pi Pa^{2}r_{+}^{2+3\omega_{q}}+9c\omega qr_{+}}.$ (46)
-To determine the winding number, we define
-$\mathcal{R}(z)=\frac{24\pi Pz^{4+3\omega_{q}}-3z^{3\omega_{q}}a^{2}+z^{2+3\omega_{q}}\left(3+8\pi Pa^{2}\right)+9cz\omega_{q}}{\tau\left(24\pi Pz^{4+3\omega_{q}}-3z^{3\omega_{q}}a^{2}+z^{2+3\omega_{q}}\left(3+8\pi Pa^{2}\right)+9cz\omega_{q}\right)+4\pi z^{3+3\omega_{q}}\left(-3+8\pi Pa^{2}\right)}.$ (47)
-For $\omega_{q}=-2/3$, this simplifies to
-$\mathcal{R}(z)=\frac{3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}}{4\pi z^{3}\left(-3+8P\pi a^{2}\right)+\left(3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}\right)\tau}.$ (48)
-Taking the denominator of the equation, we define a polynomial function $\mathcal{A}(z)$ as
-$\mathcal{A}(z)=$ $4\pi z^{3}\left(-3+8P\pi a^{2}\right)+\left(3z^{2}-6cz^{3}+24P\pi z^{4}\right.$ (49)
-$\left.+8P\pi a^{2}z^{2}-3a^{2}\right)\tau.$
-Figure 3(a) shows the variation of $r_{+}$ with $\tau$ where a generation point and an annihilation point (where $(d^{2}\tau/dr_{+}^{2})<0$) are marked with black dots.
+$\beta=\frac{4\pi r_{+}^{3+3\omega_{q}}(3-8\pi Pa^{2})}{3r_{+}^{2+3\omega_{q}}+24\pi Pr_{+}^{4+3\omega_{q}}-3a^{2}r_{+}^{3\omega_{q}}+8\pi Pa^{2}r_{+}^{2+3\omega_{q}}+9c\omega_{q}r_{+}}.$ (41)
+Figure 11 shows the variation of $r_{+}$ with $\beta$, a generation point and an annihilation point located in $\beta/r_{0}=\beta_{a}/r_{0}=3.0101$ and $\beta/r_{0}=\beta_{b}/r_{0}=4.2354$, respectively.
 These points divide the black hole into three branches: a small black hole branch, an intermediate black hole branch, and a large black hole branch.
-To compute the topological numbers, we plot the roots of $\mathcal{A}(z)$ for three cases.
-For $\tau<\tau_{a}$ or $\tau>\tau_{b}$, Figure 3(b) shows the distribution of roots when $\mathcal{A}(z)=0$.
-In these cases, there is one positive real root and one negative real root, corresponding to a winding number and topological number of 1.
-Additionally, for $\tau_{a}<\tau<\tau_{b}$.
-In these cases, there is one positive real root and one negative real root, corresponding to a winding number and topological number of 1.
-Additionally, for
-[…]
-According to the residue theorem, the winding numbers for the positive real roots are $w_{1}=1$, $w_{2}=-1$, $w_{3}=1$, resulting in the same topological number $W=1$.
-This can also be seen in Fig.
-3(c), which displays the unit vector field $n$ for $\tau_{a}<\tau<\tau_{b}$.
-Therefore, the Kerr–AdS black hole in a quintessence background exhibits a distinct topological number compared to Kerr black hole.
-In Fig.
-3(a), the generation and annihilation points are denoted by $\tau_{a}$ and $\tau_{b}$, respectively.
-The calculation of the topological number for $\tau=\tau_{a}$ using the residue theorem was discussed in the previous section and will not be repeated here.
-Below, we compute the topological number for $\tau=\tau_{b}$, where $z_{3}=z_{2}$
-[…]
-Denoting this coincident point as $z_{23}$, the characteristic function becomes
-$\mathcal{R}(z)=\frac{3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}}{(z-z_{1})(z-z_{23})^{2}(z-z_{4})}.$ (52)
-Using the residue theorem, we calculate the residue of the characteristic function at its second-order pole $z_{23}$, yielding the topological number for this point:
-$w_{23}=1.$ (53)
-(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane, a creation point and a annihilation located in $\tau/r_{0}=\tau_{a}/r_{0}=3.0101$ and $\tau/r_{0}=\tau_{b}/r_{0}=4.2354$.
-(b) The roots of $\mathcal{A}(z)$.
-(c) The unit vector field $n$ of $\tau/r_{0}=3.5$.
-Figure 3: The topological properties of the Kerr-AdS black hole surrounded by quintessence, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
-IV Einstein-Gauss-Bonnet black hole
+Figure 11: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for the Kerr-aBH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
+The characteristics of the vector field $n$ shown in Fig.
+12, where three zero points, indicated by black dots, are identified.
+These zeros partition the black hole configurations into small, intermediate, and large types.
+Fig.
+13 presents the evolution of the vector field components ($\phi^{r_{+}},\phi^{\Theta}$) traced along the closed contours $C_{i}$ depicted in Fig.
+12.
+The zero point of $\Phi_{i}$ is located at the origin.
+Contours with three zeros (from left to right) possess winding numbers +1, -1 and +1, respectively.
+The asymptotic behavior of $\phi$ along the boundary segments $I_{i}$ is summarized in Table II: at $I_{1}$ the vector field points rightward, while at $I_{3}$ it points leftward, revealing its topological number of +1.
+In the regime of low temperature ($\beta\to\infty$), the configuration gives rise to stable small black holes; conversely, when $\beta\to 0$, stable large black holes dominate the system.
+Thus, the Kerr-aBH-$\omega_{q}$ is classified within the topological class $W^{1+}$.
+Figure 12: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Kerr-aBH-$\omega_{q}$, we set $\beta/r_{0}=3.5$.
+Figure 13: Contours $\Phi_{6}$, $\Phi_{7}$ and $\Phi_{8}$ depict the variation of the vector field $\phi$ components along the pats $C_{6}$, $C_{7}$ and $C_{8}$ show in Fig.
+12 for the Kerr-aBH-$\omega_{q}$.
+VI Einstein-Gauss-Bonnet black hole surrounded by quintessence
 In this section, we focus on the thermodynamic topological number associated with black hole entropy in complex forms, such as black hole solutions surrounded by dark energy in the Einstein-Gauss-Bonnet (EGB) gravity theory.
@@
 By adjusting the state parameter $\omega_{q}$, a variety of black hole configurations can be realized.
-For example, when $\omega_{q}=0$, , corresponding to the absence of quintessence, the metric again reduces to the standard 4D EGB black hole  Glavan .
+For example, when $\omega_{q}=0$, corresponding to the absence of quintessence, the metric again reduces to the standard 4D EGB black hole Glavan .
 The function $f(r)$ and the location of the event horizon are highly sensitive to the parameters $\omega_{q}$, $q$, and the Gauss–Bonnet coupling $\alpha$.
-To ensure the physical validity of the quintessence field, $\omega_{q}$ must remain within the interval $-1<\omega_{q}<-1/3$, , which ensures a positive energy density of the dark energy component, given by  4D
+To ensure the physical validity of the quintessence field, $\omega_{q}$ must remain within the interval $1<\omega_{q}<-1/3$, , which ensures a positive energy density of the dark energy component, given by 4D
 $\rho_{q}=-6w_{q}q\frac{1}{r^{3\left(1+w_{q}\right)}}.$ (46)
-In equation (57), the "$\pm$" symbol represents the two branches of the solution.
+In equation (45), the $\pm$” symbol represents the two branches of the solution.
 For large horizon radius $r$, equation (45) reduces to the four-dimensional Schwarzschild black hole solution surrounded by quintessence matter.
@@
 $S=\frac{A}{4}[1+\frac{4}{r_{+}^{2}}\mathrm{log}(r_{+})],$ (49)
-Based on above thermodynamic quantities, the generalized off-shell free energy of the system can be derived
+here $A=4\pi r_{+}^{2}$ represents area of the black hole horizon. Based on above thermodynamic quantities, the generalized off-shell free energy of the system can be derived
 $\mathcal{F}=\frac{r_{+}}{2}(1+\frac{\alpha}{r_{+}}-\frac{2q}{r_{+}^{3\omega_{q}+1}})-\frac{\pi r_{+}^{2}(r_{+}^{2}+4\alpha\mathrm{log}(r_{+}))}{r_{+}^{2}\tau}.$ (50)
 From this, the component $\phi^{r_{+}}$ of the vector field $\phi$ can be determined
-$\phi^{r_{+}}=\frac{1}{2}+3q\omega_{q}r_{+}^{-1-3\omega_{q}}-\frac{2\pi(r_{+}^{2}+2\alpha)}{r_{+}\tau},$ (63)
+$\phi^{r_{+}}=\frac{1}{2}+3q\omega_{q}r_{+}^{-1-3\omega_{q}}-\frac{2\pi(r_{+}^{2}+2\alpha)}{r_{+}\tau}.$ (51)
 Therefore, the inverse temperature of the 4D EGB black hole is given by
-$\tau=\frac{4\pi r_{+}^{3\omega_{q}}(r_{+}^{2}+2\alpha)}{r_{+}^{1+3\omega_{q}}+6q\omega_{q}},$ (64)
-In this case, the characteristic function is:
-$\mathcal{R}(z)=$ $-\frac{z^{1+3\omega_{q}}+6q\omega_{q}}{4\pi z^{3\omega_{q}}(z^{2}+2\alpha)-\tau(z^{1+3\omega_{q}}+6q\omega_{q})}$ (65)
-$=$ $-\frac{z^{1+3\omega_{q}}+6q\omega_{q}}{\mathcal{A}(z)}.$
-Figure 4: The roots of $\mathcal{A}(z)=0$ for 4D EGB black hole. we set $\alpha=0.1$ and $q=0.1$
-We analyze the roots of $\mathcal{A}(z)=0$.
-In order to examine the poles of the complex function $\mathcal{R}(z)$ and the corresponding winding numbers, we plot $\mathcal{A}(z)$ in Fig.
-4.
-when $\tau_{c}<\tau$, the polynomial $\mathcal{A}(z)$ admits two positive real roots $z_{1}$ and $z_{2}$, satisfying $z_{1}<z_{2}$.
-Accordingly, the winding numbers are $w_{1}=1$ and $w_{2}=-1$, yielding a total topological number $W=0$.
-This result differs from the topological number of the Schwarzschild black hole under quintessence.
-Furthermore, Fig.
-5 presents the zero-point diagram of $\phi^{r_{+}}$, from which it can be seen that the 4D EGB black hole features a generation point but lacks an annihilation point.
-Figure 5: The zero points of $\phi^{r_{+}}$ in 4D and 5D EGB black holes are shown, where the black dots represent generation points. Let the parameters be $\alpha=0.1$ and $q=0.1$
-For the case of 5D, the solution of this black hole is 5DQ
+$\beta=\frac{4\pi r_{+}^{3\omega_{q}}(r_{+}^{2}+2\alpha)}{r_{+}^{1+3\omega_{q}}+6q\omega_{q}}.$ (52)
+Figure 14 presents the zero-point diagram of $\phi^{r_{+}}$, from which it can be seen that the 4D EGB black hole surrounded by quintessence (4D-EGB-BH-$\omega_{q}$) features a generation point $\beta_{c}$ but lacks an annihilation point.
+Figures 15 and 16 indicate that for $\beta>\beta_{c}$, the winding numbers of the zero points enclosed by the contours $C_{9}$ and $C_{10}$ are 1 and -1, respectively, yielding a total topological number $W=0$.
+This outcome differs from that of the Schwarzschild black hole surrounded by quintessence matter.
+Nevertheless, it can be inferred that the 4D EGB black hole and the Kerr black hole in the same matter field belong to the identical universal thermodynamic class, $W^{0+}$.
+Furthermore, the asymptotic behavior of the vector field near the boundaries $I_{i}$ is summarized in Table III.
+Figure 14: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for 4D/5D-EGB-BH-$\omega_{q}$, we set $\alpha/r_{0}^{2}=0.1$ and $qr_{0}=0.1$.
+Figure 15: The unit vector field $n$ on the ${r_{+}}-\Theta$ plane for the 4D-EGB-BH-$\omega_{q}$, we set $\beta/r_{0}=20$.
+Figure 16: Contours $\Phi_{9}$ and $\Phi_{10}$ depict the variation of the vector field $\phi$ components along the pats $C_{9}$ and $C_{10}$ show in Fig. 15 for the 4D-EGB-BH-$\omega_{q}$ case.
+For the case of 5D EGB black hole surrounded by quintessence (5D-EGB-BH-$\omega_{q}$), the solution of this black hole is 5DQ
 $f_{\pm}=1+\frac{r^{2}}{4\alpha}(1\pm\sqrt{1+\frac{8\alpha M}{r^{4}}+\frac{8\alpha q}{r^{4\omega_{q}+4}}}),$ (53)
@@
 $\tau=\frac{2\pi r_{+}^{1+4\omega_{q}}(r_{+}^{2}+4\alpha)}{r_{+}^{2+4\omega_{q}}+2q\omega_{q}}.$ (59)
-The characteristic complex function in this case is
-$\mathcal{R}(z)=$ $-\frac{z^{2+4\omega_{q}}+2q\omega_{q}}{2\pi z^{1+4\omega_{q}}(z^{2}+4\alpha)-\tau(z^{2+4\omega_{q}}+2q\omega_{q})}$ (73)
-$=$ $-\frac{z^{2+4\omega_{q}}+2q\omega_{q}}{\mathcal{A}(z)}.$
 As shown in Fig.
-5, we observe that the zeros of the vector field component $\phi^{r_{+}}$ for both 4D and 5D EGB black holes exhibit similar behavior.
-Therefore, the results for the 5D EGB black hole can be derived from those of the 4D EGB black hole.
-The findings indicate that for $\tau>\tau_{c}$, the winding numbers and the topological number are $w_{1}=1$, $w_{2}=-1$, and $W=0$, respectively.
-Table 1: In the table below, we summarize the topological numbers of the black holes studied in this work, along with their generation points (GP) and annihilation points (AP).
-We also include results reported in some references for comparison.
-Black hole spacetime Topological number GP AP
-Schwarzschild BH Wei2 -1 0 0
-Kerr BH T4 0 1 0
-Kerr-AdS BH T4 1 1 or 0 1 or 0
-Schwarzschild BH surrounded by quintessence -1 0 0
-Kerr BH surrounded by quintessence 0 1 0
-Kerr-AdS BH surrounded by quintessence 1 1 or 0 1 or 0
-4D EGB BH surrounded by quintessence 0 1 0
-5D EGB BH surrounded by quintessence 0 1 0
-V Conclusions
-This article explores the thermodynamic topological properties of black holes by utilizing the residue theorem.
-In this study, black hole solutions are regarded as defects in thermodynamic space, and a characteristic complex function is defined based on the generalized off-shell free energy.
-The singularities of this complex function are determined by the roots of the polynomial $\mathcal{A}(z)$.
-The winding number of each black hole branch is given by the residue of the characteristic function at its simple poles, which can be calculated using formula (17).
-The winding numbers reveal the local topological properties of the black hole system.
-By summing the winding numbers of all singular points, the global topological number is obtained, which aids in classifying different black hole spacetimes.
-Moreover, the winding numbers of generation points and annihilation points can be identified by the residues of the characteristic function at their second-order isolated singularities.
-We studied the winding numbers and topological numbers of a family of black holes surrounded by quintessence, with the results summarized in Table I, demonstrating that the results obtained via the residue method agree with those calculated through the topological current theory.
-Within the chosen black hole models, we find that the Schwarzschild black hole in quintessence has a topological number of -1, with no generation or annihilation points.
-In contrast, the Kerr black hole and the 4D/5D Einstein-Gauss-Bonnet black holes under quintessence background have a topological number of 0, each featuring one generation point but no annihilation point.
-For the Kerr-AdS black hole in quintessence, its topological number matches that of the Kerr-AdS black hole without quintessence, with one generation point and one annihilation point identified.
-These results indicate that both the rotation parameter and the Gauss-Bonnet coupling parameter under the quintessence background alter the topological class of black holes.
+14, we observe that the zeros of the vector field component $\phi^{r_{+}}$ for both 4D and 5D EGB black holes exhibit similar behavior.
+The results for the 5D EGB black hole can be derived from those of the 4D EGB black hole.
+Therefore, the 5D-EGB-BH-$\omega_{q}$ belongs to the specific topological class, $W^{0-}$.
+Table 3: The orientation is indicated by the arrow of $\phi^{r_{+}}$ for the 4D/5D-EGB-BH-$\omega_{q}$ and the topological numbers..
+BH spacetime $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
+4D-EGB-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\leftarrow$ $\downarrow$ 0
+5D-EGB-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\leftarrow$ $\downarrow$ 0
+Table 4: In the table below, we summarize the universal thermodynamic topological classifications of the black holes surrounded by quintessence, together with a summary of their corresponding thermodynamic behaviors and their generation points (GP) and annihilation points (AP).
+Black hole spacetime W Classes Innermost Outermos Low T High T DP
+Sch-BH T14 $W^{1-}$ unstable unstable unstable large unstable small in pairs
+Kerr-BH T17 $W^{0+}$ stable unstable unstable large + stable small no one more GP
+Sch-aBH T14 $W^{0-}$ unstable stable no unstable small + stable large one more AP
+Sch-BH-$\omega_{q}$ $W^{1-}$ unstable unstable unstable large unstable small in pairs
+Kerr-BH-$\omega_{q}$ $W^{0+}$ stable unstable stable small + unstable large no one more GP
+Sch-aBH-$\omega_{q}$ $W^{0-}$ unstable stable no unstable small + stable large one more AP
+Kerr-aBH-$\omega_{q}$ $W^{1+}$ stable stable stable small stable large in pair
+4D-EGB-BH-$\omega_{q}$ $W^{0+}$ stable unstable unstable large + stable small no one more GP
+5D-EGB-BH-$\omega_{q}$ $W^{0+}$ stable unstable unstable large + stable small no one more GP
+VII Conclusions
+In this work, we have systematically investigated the universal thermodynamic topological classifications of a family of black holes surrounded by quintessence, including Schwarzschild, Kerr, AdS, and Einstein–Gauss–Bonnet (EGB) spacetimes.
+The results are summarized in Table IV.
+Our analysis reveals that the Schwarzschild black hole in a quintessence background belongs to the $W^{1-}$ class, showing thermodynamic instability in both low- and high-temperature regimes.
+The introduction of rotation fundamentally modifies the topological characteristics: the Kerr black hole transitions to the $W^{0+}$ class, allowing coexistence of a stable small and an unstable large black hole at low temperatures, while no stable configurations persist at high temperatures.
+The presence of a quintessence field does not alter this universal classification.
+For the AdS counterparts, the thermodynamic topology becomes markedly distinct.
+The Schwarzschild–AdS black hole corresponds to the $W^{0-}$ class, featuring the emergence of an unstable small and a stable large black hole only at high temperatures, with no black hole solution at low temperatures.
+In contrast, the Kerr–AdS configuration is classified as $W^{1+}$, where both small and large black holes can maintain stability across temperature ranges.
+This distinction highlights the significant influence of the AdS curvature and rotational degrees of freedom on the thermodynamic topology.
+Furthermore, the inclusion of higher-order curvature corrections through the Gauss–Bonnet term in both 4D and 5D spacetimes preserves the $W^{0+}$ classification.
+These cases exhibit similar thermodynamic structures to the Kerr black hole in quintessence, characterized by the coexistence of stable and unstable branches at low temperatures and the disappearance of black hole states at high temperatures.
+Overall, the comparative analysis demonstrates that the topological classification of black holes is sensitive to intrinsic parameters such as rotation and higher-curvature couplings, while the quintessence field mainly shifts the temperature-dependent stability ranges without altering the topological class itself.
+This work thus provides a new perspective for understanding the global stability structures of black holes in dark energy backgrounds.
 Acknowledgments
-The research was partially supported by the National Natural Science Foundation of China (Grant No.12265007) and the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07, QJJ-[2022]-314).
+We are very grateful to Dr.
+P.
+K.
+Yerra for helpful discussion.
+We also acknowledge the anonymous referees for their valuable comments on improving our paper.
+The research was partially supported by the National Natural Science Foundation of China (Grant No.12265007).
 Also, the research of H.H.